OpenStudy (anonymous):

Solve for x: 1 over 4 (2x − 14) = 4

OpenStudy (anonymous):

(2x - 14) / 4 = 4 2x - 14 = 4(4) = 16 2x = 16 + 14 = 30 x = 30 / 2 = 15

OpenStudy (anonymous):

So x= 15

OpenStudy (anonymous):

so x = 15

OpenStudy (ybarrap):

$$ \Large \bf \dfrac{1}{4(2x-14)}=4\\\\ \Large \bf \dfrac{2x-14}{4(2x-14)}=4(2x-14)\\\\ \Large \bf \dfrac{1}{4}=4(2x-14)\\\\ \Large \bf \dfrac{1}{4(4)}=\frac{4(2x-14)}{4}\\\\ \Large \bf \dfrac{1}{16}=2x-14\\\\ \Large \bf \dfrac{1}{16}+14=2x-14+14\\\\ \Large \bf \dfrac{1}{16}+14=2x\\\\ \Large \bf 2x=\dfrac{1}{16}+14\\\\ \Large \bf \frac{2x}{2}=\dfrac{\dfrac{1}{16}+14}{2}\\\\ \Large \bf x=\dfrac{1}{32}+7\\\\ \Large\text{Answer: }\Large \bf x=7\dfrac{1}{32}\\\\ $$